Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
Off-Campus Internet Connection Apr 27 2011 - 06:00 Users connecting to off campus services We are currently experiencing data packet loss through the CENIC network. TNS is working closely with CENIC to resolve issue
Off-Campus Internet Connection Apr 23 2011 - 05:00 User connecting to off campus services We are currently experiencing data packet loss through the CENIC network. TNS is working closely with CENIC to resolve issue
Other Apr 21 2011 - 14:24 All Services located in the HSU Data Center HVAC failure. The problem has been fixed, the temperature is back to normal.
KBOX Ticketing System Apr 18 2011 - 13:30 Anyone using the Help Desk Ticketing System The problem appears to be caused by some looping tickets. The tickets have been cleaned up and a rule has been put in place to prevent this problem in the future.
Campus Data Network Apr 17 2011 - 08:00 Users inside Theater Arts (TA) & Student Health Center (SHC) only In December 2010, a secondary L2 up-link was installed to TA and SHC as part of the ITRP2 Refresh Project. In order to be fully redundant, the L3 interfaces on two core switches must be re-configured to support Hot Standby Route Protocol (HSRP). While these changes are being made, connectivity will be briefly interrupted. This must be completed prior to final ITRP2 project fail-over testing on April 30th.
Campus Data Network Apr 15 2011 - 07:00 Potentially affecting entire campus data network (wired & wireless), however the risk is less than 1%. ***UPDATE APR152011: The standby supervisor was successfully installed without any impact to the network.*** Original Message: In early March, network technicians removed a standby supervisor from the core network switch in Siemens Hall. The standby supervisor was to be used in SBS if a software upgrade did not resolve instability issues there. Fortunately, the SBS issues were resolved and the standby supervisor must now be re-installed in SH. NO IMPACT IS EXPECTED and this is being posted only as a precaution.
Hyperion Apr 17 2011 (All day) All users accessing Data Warehouse and Reporting Services (Hyperion, OBI, etc) Monthly ITS Maintenance Window
Other Apr 17 2011 - 05:00 All Users Monthly ITS Maintenance Window
CMS/Peoplesoft Apr 13 2011 - 10:59 UPDATE: PeopleSoft is up.
CMS/Peoplesoft Apr 13 2011 - 14:37 PeopleSoft is up.
CMS/Peoplesoft Apr 15 2011 - 21:00 Anyone who uses PeopleSoft CFS (Finance) CMS Central (Unisys Data Center) will upgrade the production database to Solaris 10 and Oracle CRS. The PeopleSoft application and data will not be affected, just unavailable during this maintenance effort.
CMS/Peoplesoft Apr 7 2011 - 18:00 PeopleSoft CFS (Finance) CMS Central is applying changes to the software. The changes do not affect the features and functionality used by Finance or Campus users.
Pharos Pay-for-Print Apr 10 2011 - 07:00 Students printing in computer labs The University Center (which hosts the CBORD system) will be performing electrical repairs, which require the CBORD system to be powered down. The Pharos system relies on CBORD to charge C-cards for printing.
Network Folders Apr 5 2011 (All day) All Folders Users The Winbind service on server Users3 became stale, and had to be restarted.
Other Apr 4 2011 - 21:00 Very few users Virtual servers are being moved to a different datastore
Campus Data Network Mar 10 2011 - 11:00 Many Student Business Services network users UPDATE:**** Problem has been solved and there have been no new issues. We apologize for any inconvenience this may have caused. ****
Other Apr 1 2011 - 10:17 Data Warehouse Users ****UPDATE****: Data Warehouse is once again available. Check back for updates.
CMS/Peoplesoft Mar 30 2011 - 18:00 Anyone who uses PeopleSoft CFS (Finance) CMS Central is applying projects for Third Wave campuses who are going live.
Off-Campus Internet Connection Mar 28 2011 - 15:45 Primary link to Sunnyvale down - secondary link up so campus has internet. No one at this time (have secondary circuit). Just notice for information. At this time, AT&T is investigating
CMS/Peoplesoft Mar 25 2011 - 21:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions CMS Central (Unisys Data Center) database upgrade to Solaris 10 and Oracle CRS.
Account Center Mar 23 2011 - 10:23 Users trying to access Account Center UPDATE: *****Account Center is once again available.***** Account Center has been restarted.
Computer Labs Mar 22 2011 - 11:00 Users in Smart Classrooms & Interdisciplinary Labs Problem is due to a KBox glitch. The vendor has been contacted and is working on the issue.
KBOX Ticketing System Mar 21 2011 - 10:29 KBox users UPDATE: **** KBOX has been restarted.**** KBox ticketing system is going to be restarted. False tickets will be deleted. This should speed up KBox.
Off-Campus Internet Connection Mar 18 2011 - 16:00 Potential Risk of Losing Connection to Internet Earlier this month, the University's secondary Internet connection was upgraded from 155MMbps to 1Gbps. Unfortunately, resources to test the fail-over from primary to secondary were not available. HSU Network Technicians will perform a simulated failure of the Primary Internet connection to verify the Secondary link is fully functional.
Account Center Mar 18 2011 - 09:59 Users trying to reset their password UPDATE: Password reset is available. We are checking into the problem.
CMS/Peoplesoft Mar 16 2011 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions The Orientation modification (AD), Checklist Fix (AD), and DARS Fix will be installed.
Account Center Mar 15 2011 - 08:41 Users trying to change their password We are checking into the problem. Check back for update.
Campus Data Network Mar 14 2011 - 11:47 Student Business Services building Data Services were restored by 1:12pm. TNS carried out complete hardware level tests on all switch modules after we were notified of a connectivity failure at 11:47am. All hardware diagnostics passed which suggests a software (IOS) bug. The latest IOS software image was loaded onto this switch by 1pm and the switch was returned to full service and connectivity restored by 1:12pm. We are hopeful this will resolve the situation permanently, however we have started planning a full switch replacement should it be needed.
Other Mar 14 2011 - 14:40 Network Folders/User Web Sites/ NRS-Labs server access The emergency generator failed resulting in loss of power to some systems including a storage array (SAN). When the SAN was returned to operation numerous systems were operating in read-only mode. There were numerous hardware failures, and ldap, shibboleth IdP, Network Folders, NRS-Labs, users, user-web were affected.
Other Mar 14 2011 - 13:57 UPDATE: Resource 25 users UPDATE: Resource 25 is expected to be down all day.
feedback